# Bayesian mixture modeling using a mixture of finite mixtures with normalized inverse Gaussian weights

### v1

@@ 3.4.2 Network analysis
As an application of the proposed method, we perform community detection on network data. Community detection is the task of identifying dense subclasses in network data and corresponds to clustering and estimating the number of components, called the number of communities in network analysis. Note that the number of components of finite mixture models is equivalent to the number of communities in the network, and both are denoted $M$. Estimating the number of communities is an important problem and various methods have been proposed (Shi and Malik, 2000; Newman, 2004; White and Smyth, 2005). From a model-based perspective, it is essentially the same as estimating the number of components in a finite-mixture model, and we can apply MFM. The stochastic block model is a famous statistical model of network data (Henze, 1986; Nowicki and Snijders, 2001; Geng et al., 2019), which assumes a stochastic block structure behind and specifies the community structure by estimating the probability of edges being drawn between each group. We estimate the number of communities with the MFM in the framework of this stochastic block model. Geng et al. (2019) proposed a stochastic block model based on MFM with Dirichlet weights, and also constructed a similar algorithm to Miller and Harrison (2018).
MFM can be easily applied to community detection by modifying the kernel. Data $y$ are replaced by the adjacency matrix $A\in\mathbb{R}^{n\times n}$, where $A=(A_{ij})\in\{0,1\}^{n\times n}$ and $n$ is the size of the node. When $A_{ij}=1$, this indicates that an edge is drawn from the $i$ th node to the $j$ th node, and when $A_{ij}=0$, this does not. For simplicity, we assume that the adjacency matrix is not direct and does not have a self-loop, in that $A_{i,j}=A_{ji}$ and $A_{ii}=0$, where $1\leq i<j\leq n$. The stochastic block model is formulated as follows,
$\begin{split}&A_{ij}|Q,M\sim\text{Bernoulli}(\theta_{ij}),\quad\theta_{ij}=Q_{c_{i}c_{j}},\quad 1\leq i<j\leq n,\\ &Q_{rs}\sim\text{Beta}(a_{Q},b_{Q}),\quad 1\leq r\leq s\leq M,\end{split}$ (3.5)
where $a_{Q},\ b_{Q}>0$ and $M$ is the number of communities. $Q=(Q_{rs})\in[0,1]^{M\times M}$ is a symmetric matrix and defines the stochastic block structure of a network. Each element $Q_{rs}$ represents the probability that an edge is drawn between any node belonging to the community label $r$ and any node belonging to the community label $s$. To perform community detection based on the proposed model, we just set the Bernoulli likelihood as the kernel.
@@ 3.5 Evaluation of the number of empty components
From the point of view of the interpretability of the model, the generalization performance, and the computational efficiency, it is reasonable that the number $M_{na}$ should be small. For the full conditional distribution of $M_{na}$, the following inequality holds, where $M-1|\Lambda\sim\mathrm{Poisson}(\Lambda)\ \text{and}\ \Lambda\sim\mathrm{Gamma}(a_{\Lambda},b_{\Lambda})\ \text{for}\ a_{\Lambda},b_{\Lambda}>0$.
@@
$\mathbb{P}(M_{na}\geq 1|U_{n}=u,\ k,\ M=m,\ \Lambda)$ $\leq\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right),$ (3.6)
$\mathbb{P}(M_{na}\geq 1|U_{n}=u,\ k,\ M=m)$ $\leq\psi(u)\frac{a_{\Lambda}}{b_{\Lambda}}\left(1+\frac{1}{k}\right)$ (3.7)
holds, where $\psi$ is the Laplace transform of the probability distribution for $h$.
Proof.
When $m\neq 0$,
$\mathbb{P}(M_{na}=m|U_{n}=u,\ k,\ M=m,\ \Lambda)=\frac{k}{\Lambda\psi(u)+k}\mathcal{P}_{0}(m;\Lambda\psi(u))+\frac{\Lambda\psi(u)}{\Lambda\psi(u)+k}\mathcal{P}_{1}(m;\Lambda\psi(u)).$
Thus, the conditional expectation of $M_{na}$ is
$\mathbb{E}[M_{na}|U_{n}=u,\ k,\ M=m,\ \Lambda]$ $=\frac{k}{\Lambda\psi(u)+k}\times\Lambda\psi(u)+\frac{\Lambda\psi(u)}{\Lambda\psi(u)+k}\times(\Lambda\psi(u)+1)$
$=\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right).$
Furthermore,
$\mathbb{E}[M_{na}|U_{n}=u,\ k,\ M=m]$ $=\mathbb{E}_{\Lambda}[\mathbb{E}[M_{na}|U_{n}=u,\ k,\ M=m,\ \Lambda]]$
Finally, the result follows from Markov’s inequality. ∎
[…]
$\psi_{\text{Inv-Ga}}(u)=\exp\left(\alpha\left(1-\sqrt{1+2u}\right)\right),\quad u\geq 0,$ (3.8)
$\psi_{\text{Ga}}(u)=\left(\frac{\eta}{\eta+u}\right)^{\gamma},\quad u\geq 0.$ (3.9)
Laplace transforms $\psi_{\text{Inv-Ga}}$ and $\psi_{\text{Ga}}$ are decreasing functions with respect to $u$. The former has exponential decay, while the latter is polynomial. Figure 1 shows the graphs of the Laplace transforms of inverse Gaussian and gamma distributions when the shape parameters are $\alpha,\gamma=1.0,0.2,10^{-1},10^{-2},10^{-3}$ and the scale parameters are $1$. It can be seen that $\psi_{\text{Inv-Ga}}$ decreases much faster than $\psi_{\text{Ga}}$. The speed of this decrease has a significant impact on the appearance of the empty component in estimating the number of components.
[…]
@@ 4 Empirical demonstrations
We evaluate the performance of the MFM-Inv-Ga and MFM-Ga methods through some numerical experiments. Recall that $\gamma$ is the shape parameter of the prior distribution of unnormalized weights in the proposed method (MFM-Inv-Ga), while $\alpha$ is that of MFM-Ga.
4.1 Inference for the number of mixture components and clustering
In this subsection, we illustrate the performance of clustering and inference for the number of components using artificial and real data.
@@ 4.1.1 Artificial data
In this simulation, we assume that $M_{\mathrm{true}}=3$. We generate data from the following multivariate normal distribution:
$f(y|\mu_{1},\mu_{2},\mu_{3})=0.8\cdot N_{2}(y|\mu_{1},I_{2})+0.1\cdot N_{2}(y|\mu_{2},I_{2})+0.1\cdot N_{2}(y|\mu_{3},I_{2}),$
[…]
  • The posterior mean of the number of components $M$.
  • The posterior mean of the rand index. The rand index is a measure of the clustering fitting, and the value takes $[0,1]$. When it is close to $1$, the assignment estimate is reasonable.
  • The posterior probability that the number of empty components $M_{na}$ is equal to 00.
The respective averages over $50$ repetitions are denoted by $\widehat{\mathrm{RI}}$ and ^$\widehat{\mathbb{P}}(M_{na}=0)$.
We report the posterior mean of $M$ in Figure 2. It is observed that the results of the MFM-Inv-Ga method remain almost the same even if the shape parameter is varied. However, the MFM-Ga method tends to overestimate the number of mixture components as the shape parameter decreases.
Figure 2: Box plots of the posterior mean of the number of components $M$ for each shape parameter. From left to right, shape parameters are set as $\gamma,\alpha=1,0.2,10^{-1},10^{-2},10^{-3}$.
Table 1 shows the results of the clustering performance and the posterior probability that $M_{na}$ is equal to 00. For clustering, the MFM-Inv-Ga and MFM-Ga methods are comparable. Both methods have reasonable clustering accuracy, only slightly better for MFM-Ga than for MFM-Inv-Ga. As the shape parameters $\alpha$ and $\gamma$ decrease, $\widehat{\mathrm{RI}}_{i}$ increases slightly and its standard deviation decreases. This is natural given that for data that have components with relatively very large cluster sizes, a suitable prior on a simplex is one with large mass at the edges or vertices. It is important to note that Figure 2 and Table 1 show that the MFM-Inv-Ga method produces few empty components for all scenarios and provides reasonable estimates of $M$, but not MFM-Ga.
Table 1: Results of the rand index $\widehat{\mathrm{RI}}_{i}$ and posterior probability that the number of empty components is equal to zero ^$\widehat{\mathbb{P}}(M_{na}=0)$ (their standard deviations are shown in parentheses) averaged over 50 Monte Carlo replications for each shape parameters.
$\widehat{\mathrm{RI}}_{i}$
$\alpha$ and $\gamma$ $1$ $0.2$ $10^{-1}$ $10^{-2}$ $10^{-3}$
MFM-Inv-Ga $0.993$ $0.994$ $0.995$ $0.995$ $0.995$
$(0.013)$ $(0.006)$ $(0.006)$ $(0.004)$ $(0.005)$
MFM-Ga $0.995$ $0.995$ $0.995$ $0.996$ $0.996$
$(0.006)$ $(0.004)$ $(0.003)$ $(0.004)$ $(0.003)$
^$\widehat{\mathbb{P}}(M_{na}=0)$
$\alpha$ and $\gamma$ $1$ $0.2$ $10^{-1}$ $10^{-2}$ $10^{-3}$
MFM-Inv-Ga $1.000$ $0.995$ $0.982$ $0.947$ $0.952$
$(0.001)$ $(0.008)$ $(0.032)$ $(0.135)$ $(0.087)$
MFM-Ga $0.979$ $0.555$ $0.328$ $0.088$ $0.066$
$(0.008)$ $(0.080)$ $(0.066)$ $(0.012)$ $(0.013)$
This can be seen in Table 2. It can also be seen that MFM-Inv-Ga assigns a very high posterior probability to $M_{\mathrm{true}}$ than MFM-Ga, and the behavior of the posterior distributions for $k$ and $M$ is similar, together with Table 1.
MFM-Inv-Ga
$M=1$ $M=2$ $M=3$ $M=4$ $M=5$ $M\geq 6$
$\alpha=1$ $0.000$ $0.000$ $\bf{0.973}$ $0.027$ $0.000$ $0.000$
$(0.000)$ $(0.000)$ $(0.093)$ $(0.093)$ $(0.000)$ $(0.000)$
$\alpha=0.2$ $0.000$ $0.080$ $\bf{0.886}$ $0.033$ $0.001$ $0.000$
$(0.000)$ $(0.274)$ $(0.268)$ $(0.048)$ $(0.003)$ $(0.000)$
$\alpha=10^{-1}$ $0.000$ $0.000$ $\bf{0.942}$ $0.051$ $0.006$ $0.001$
$(0.000)$ $(0.000)$ $(0.084)$ $(0.069)$ $(0.013)$ $(0.005)$
$\alpha=10^{-2}$ $0.000$ $0.020$ $\bf{0.890}$ $0.063$ $0.015$ $0.012$
$(0.000)$ $(0.141)$ $(0.199)$ $(0.078)$ $(0.039)$ $(0.048)$
$\alpha=10^{-3}$ $0.000$ $0.050$ $\bf{0.860}$ $0.068$ $0.015$ $0.007$
$(0.000)$ $(0.209)$ $(0.225)$ $(0.084)$ $(0.029)$ $(0.020)$
MFM-Ga
$\gamma=1$ $0.000$ $0.000$ $\bf{0.948}$ $0.049$ $0.002$ $0.000$
$(0.000)$ $(0.000)$ $(0.055)$ $(0.049)$ $(0.006)$ $(0.001)$
$\gamma=0.2$ $0.000$ $0.010$ $\bf{0.508}$ $0.301$ $0.121$ $0.061$
$(0.000)$ $(0.070)$ $(0.093)$ $(0.037)$ $(0.037)$ $(0.035)$
$\gamma=10^{-1}$ $0.000$ $0.000$ $\bf{0.307}$ $0.306$ $0.196$ $0.191$
$(0.000)$ $(0.007)$ $(0.093)$ $(0.037)$ $(0.037)$ $(0.035)$
$\gamma=10^{-2}$ $0.000$ $0.003$ $0.088$ $0.155$ $\bf{0.176}$ $0.578$
$(0.000)$ $(0.002)$ $(0.021)$ $(0.016)$ $(0.013)$ $(0.071)$
$\gamma=10^{-3}$ $0.000$ $0.003$ $0.066$ $0.132$ $\bf{0.163}$ $0.636$
$(0.000)$ $(0.020)$ $(0.019)$ $(0.013)$ $(0.011)$ $(0.072)$
Table 3 shows the CPU times of MFM-Inv-Ga and MFM-Ga for $50$ repetitions of MCMC with $2000$ iterations, where $\alpha,\gamma=1.0$ and $10^{-3}$. When $\alpha=\gamma=1.0$, MFM-Inv-Ga is faster than MFM-Ga on average, but MFM-Inv-Ga has more variability than MFM-Ga. On the other hand, MFM-Ga is very time-consuming in $\gamma=10^{-3}$, because many empty components are created. As a result, MFM-Inv-Ga has the same clustering accuracy as MFM-Ga and outperforms MFM-Ga in terms of $M$ estimation and CPU time.
Table 3: Comparison of CPU times (in seconds) and their standard deviations (shown in parentheses) for artificial data averaged over $50$ replications, $\alpha,\ \gamma=1.0,10^{-3}$
MFM-Inv-Ga MFM-Ga
$\alpha=\gamma=1$ $88.545$ $90.898$
$(5.686)$ $(2.774)$
$\alpha=\gamma=10^{-3}$ $87.061$ $192.234$
$(12.286)$ $(8.932)$

### v2

@@ 3.4.2 Network analysis
As an application of the proposed method, we perform community detection on network data. Community detection is the task of identifying dense subclasses in network data, corresponds to clustering and estimating the number of components, called the number of communities in network analysis. Note that the number of components of finite mixture models is equivalent to the number of communities in the network, and both are denoted $M$. Estimating the number of communities is an important problem and various methods have been proposed (shi2000normalized; newman2004detecting; white2005spectral). From a model-based perspective, it is essentially the same as estimating the number of components in a finite mixture model, and we can apply MFM. The stochastic block model is a famous statistical model of network data (holland1983stochastic), which assumes a stochastic block structure behind and specifies the community structure by estimating the probability of edges being drawn between each group. geng2019probabilistic proposed a stochastic block model based on MFM with Dirichlet weights, and also constructed a similar algorithm to miller2018mixture.
MFM can be easily applied to community detection by modifying the kernel. Data $y$ are replaced by the adjacency matrix $A\in\mathbb{R}^{n\times n}$, where $A=(A_{ij})\in\{0,1\}^{n\times n}$ and $n$ is the number of the node. When $A_{ij}=1$, this indicates that an edge is drawn from the $i$ th node to the $j$ th node, and when $A_{ij}=0$, this does not. For simplicity, we assume that the adjacency matrix is not directed and does not have a self-loop, in that $A_{ij}=A_{ji}$ and $A_{ii}=0$, where $1\leq i<j\leq n$. The stochastic block model is formulated as follows,
$\begin{split}&A_{ij}\mid Q,M\sim\text{Bernoulli}(\theta_{ij}),\quad\theta_{ij}=Q_{c_{i}c_{j}},\quad 1\leq i<j\leq n,\\ &Q_{rs}\sim\text{Beta}(a_{Q},b_{Q}),\quad 1\leq r\leq s\leq M,\end{split}$ (3.7)
where $a_{Q},\ b_{Q}>0$ and $M$ is the number of communities. $Q=(Q_{rs})\in[0,1]^{M\times M}$ is a symmetric matrix and defines the stochastic block structure of a network. Each element $Q_{rs}$ represents the probability that an edge is drawn between any node belonging to the community label $r$ and any node belonging to the community label $s$. To perform community detection for the proposed model, we just set the Bernoulli likelihood as the kernel.
@@ 3.5 Evaluation of the number of empty components
For the full conditional distribution of $M_{na}$, the following inequality holds, where $M-1\mid\Lambda\sim\mathrm{Poisson}(\Lambda)\ \text{and}\ \Lambda\sim\mathrm{Gamma}(a_{\Lambda},b_{\Lambda})\ \text{for}\ a_{\Lambda},b_{\Lambda}>0$.
@@
$\mathbb{P}(M_{na}\geq 1\mid U_{n}=u,\ k,\ \Lambda)$ $\leq\Lambda\psi(u)\left(1+\frac{1}{\Lambda\psi(u)+k}\right),$ (3.8)
$\mathbb{P}(M_{na}\geq 1\mid U_{n}=u,\ k)$ $\leq\psi(u)\frac{a_{\Lambda}}{b_{\Lambda}}\left(1+\frac{1}{k}\right)$ (3.9)
hold, where $\psi$ is the Laplace transform of the density $h$.
The proof of the proposition is given in the Supplementary Material. Proposition 3.1 shows that $\psi(u)$ plays an important role in generating empty components. The critical difference between the inverse Gaussian and gamma distributions is the Laplace transform. Let $\psi_{\text{Gamma}}(u)$ be the Laplace transforms of $\text{Gamma}(\gamma,1)$:
$\psi_{\text{Gamma}}(u)=\left(\frac{1}{1+u}\right)^{\gamma},\quad u\geq 0.$ (3.10)
Laplace transforms $\psi_{\text{IGau}}$ and $\psi_{\text{Gamma}}$ are decreasing functions with respect to $u$. The former has exponential decay, while the latter is polynomial. Figure 1 shows the graphs of the Laplace transforms of inverse Gaussian and gamma distributions when the shape parameters are $\alpha,\gamma=1.0,0.2,10^{-1},10^{-2},10^{-3}$ and the scale parameters are $1$. It can be seen that $\psi_{\text{IGau}}$ decreases much faster than $\psi_{\text{Gamma}}$, contributing to the suppression of $M_{na}$.
Figure 1: When shape parameters are $\alpha=\gamma=1,0.2,10^{-1},10^{-2},10^{-3}$, left graphs are Laplace transforms of $\text{IGau}(\alpha,1)$, right graphs are Laplace transforms of $\text{Gamma}(\gamma,1)$.
@@
3.6 Relationship with the telescoping sampling
In this section, we clarify the connection between the blocked Gibbs sampling and the telescoping sampling. The blocked Gibbs sampling is derived from the posterior distribution of a random measure induced by the Norm-IFPP, whereas the telescoping sampling is a partially marginalized algorithm constructed from the EPPF. Although the two methods arise from fundamentally different formulations, they share several similarities, including the sampling of $M$. To the best of our knowledge, their relationship has not been clarified so far, as both methods were proposed independently around the same time.
First, we explain that the blocked Gibbs sampling can be derived in a manner similar to the telescoping sampling scheme. Specifically, it can be derived from both the full mixture posterior $p(M,S,U_{n},\tau_{1:M},c\mid y)$ and the collapsed mixture posterior $p(M,\mathcal{C},U_{n},\tau_{1:k}\mid y)$, where $\mathcal{C}=\{\mathcal{C}_{1},\dots,\mathcal{C}_{k}\}$ is a partition of $\{1,\dots,n\}$ into $k$ blocks, where $\mid\mathcal{C}_{i}\mid=n_{i}$ for $i=1,\dots,k$ and $\sum_{i=1}^{k}n_{i}=n$, and the hyperparameters $\rho$ and $\eta$ are omitted for simplicity. This full mixture posterior is used to sample $\tau_{1:M}$, $c$, $S$ and $U_{n}$ (Step 1, 2, 5, 6, 7 in Algorithm 1). The collapsed mixture posterior is obtained by marginalizing out $S$, the unallocated parameters, and all assignments $c$ that induce the same partition $\mathcal{C}$ in the full mixture posterior. Thus, by Theorem 4.1 in argiento2022infinity, the conditional distribution of $M$ given $U_{n}=u,\mathcal{C}$ is
$\mathbb{P}(M=m\mid\mathcal{C},U_{n}=u)\propto\frac{m!}{(m-k)!}\psi(u)^{m-k}q_{M}(m),\quad m=k,k+1,\dots.$
This full conditional distribution of $M$ is equivalent to the equation (2.7), shifted by $k$.
The latent variable $U_{n}$ is essential distinction between the two methods, while the blocked Gibbs sampling shares many advantages with the telescoping sampling discussed in fruhwirth2021generalized. In other words, the blocked Gibbs sampling can be regarded as a telescoping sampling augmented with the latent variable $U_{n}$. By this latent variable, the full conditional distribution $M_{na}$ (or $M$) can be easily sampled when $q_{M}$ is Poisson distribution or negative binomial distribution. This is a major advantage over the telescoping sampling, because, in practice, an upper bound $M_{\mathrm{max}}$ is required and $M$ is sampled from a multinomial distribution over $\{k,\dots,M_{\mathrm{max}}\}$ due to the complex form of $M$’s full conditional distribution.
3.7 Extended telescoping sampling
In this section, we extend the telescoping sampling so that it can be applied to model (2.3), thereby relaxing the Dirichlet restriction in fruhwirth2021generalized.
The conditional distributions for $\mathcal{C}$, given $M=m$ and $U_{n}=u$, and for $M$, given $\mathcal{C},U_{n}=u$ are required to construct our algorithm under (2.3).
Proposition 3.2.
Let $\mathcal{C}=\{\mathcal{C}_{1},\dots,\mathcal{C}_{k}\}$ be a set partition with $n_{j}=|\mathcal{C}_{j}|$ and $\sum_{j=1}^{k}n_{j}=n$. Under (2.3), the EPPF is the following:
$p(\mathcal{C})=\int_{0}^{\infty}\frac{u^{n-1}}{\Gamma(n)}\sum_{m=k}^{\infty}\frac{m!}{(m-k)!}\psi(u;m)^{m-k}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}q_{M}(m)\mathrm{d}u,\quad$ (3.11)
where $\psi(u;m)$ and $\kappa(u;n,m)$ are the Laplace transform and cumulant of $h_{M}$, respectively. Moreover, the conditional probability of $M=m$, for $m=k,k+1,\dots,$ is proportional to the following:
$\mathbb{P}(M=m\mid U_{n}=u,\mathcal{C})\propto\frac{m!}{(m-k)!}\psi(u;m)^{m-k}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}q_{M}(m).$ (3.12)
From Proposition 3.2, the Algorithm 2 is immediately derived, showing only the parts that differ from Algorithm 1. If $h_{M}$ is $\text{Gamma}(\gamma/M,1)$ or $\text{IGau}(\alpha/M,1)$, Algorithm 2 can be applied to the DMFM based on the Dirichlet and normalized inverse Gaussian distributions, respectively. The main difference from Algorithm 1 is that the Laplace transform and the cumulant of $h_{M}$ depend on $M$. Since the full conditional distributions of $M$ (or $M_{na}$) and the summation $\sum_{m=k}^{\infty}\frac{m!}{(m-k)!}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}\psi(u;n_{j},m)^{m-k}q_{M}(m)$ strongly depend on $\psi(u;m)$ and $\kappa(u;n,m)$, they cannot, in general, be expressed in a simple form. Thus, Step 3 in Algorithm 2 is replaced by sampling from the full mixture posterior, and $M$ is not marginalized out in Step 3’, as in fruhwirth2021generalized. In addition, in Step 4, it is practically necessary to set an upper bound $M_{\text{max}}$ and sample from a multinomial distribution.
Algorithm 2 Extended telescoping sampling
Step 3. Sample the hyperparameter $\eta$ of the $q_{M}$ from
$p(\eta\mid\mathrm{rest})\propto p(\eta)q_{M}(m).$
Step 3’. Sample the hyperparameter $\rho$ of the $h$ from
$p(\rho\mid\mathrm{rest})\propto\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}\psi(u;m)^{m-k}p(\rho).$
Step 4. Sample $M$ from the following discrete distribution
$\mathbb{P}(M=m\mid\mathrm{rest})\propto\frac{m!}{(m-k)!}\left\{\prod_{j=1}^{k}\kappa(u;n_{j},m)\right\}\psi(u;m)^{m-k}q_{M}(m),\quad m=k,k+1,\dots.$
@@ 4 Empirical demonstrations
We evaluate the performance of the MFM-IGau and MFM-Ga methods through some numerical experiments. Recall that $\gamma$ is the shape parameter of the prior distribution of unnormalized weights in the proposed method (MFM-IGau), while $\alpha$ is that of MFM-Ga.
4.1 Clustering performance
In this subsection, we compare the clustering performance of the proposed methods, MFM-IGau and DMFM-IGau, with MFM-Ga and DMFM-Ga, using artificial and real data.
@@ 4.1.1 Artificial data
In this simulation, we assume that $M_{\mathrm{true}}=k_{\mathrm{true}}=8$, and data are generated from the following multivariate normal distribution mixture:
$f(y\mid\mu_{1},\dots,\mu_{8})=\sum_{i=1}^{8}w_{i}N_{2}(y\mid\mu_{i},\sigma^{2}I_{2}),$
where $w_{i}=i^{-1.75}/\sum_{i=1}^{8}w_{i}$, $\mu_{1}=(2,0)^{\top},\mu_{2}=(2,5)^{\top},\mu_{3}=(6,0)^{\top},\mu_{4}=(6,5)^{\top},\mu_{5}=(10,0)^{\top},\mu_{6}=(10,5)^{\top},\mu_{7}=(14,0)^{\top},\mu_{8}=(14,5)^{\top}$, $\sigma^{2}=0.45$ and $I_{2}$ is the $2\times 2$ identity matrix. We set $n=50,100,200,300,400$ and generate $50$ datasets for each $n$.
This data generating process represents a situation in which many mixture components have small weights, so that not all components necessarily contribute to the observed data. When $n$ is small, it is difficult to obtain observations from all components, whereas as $n$ increases, it becomes more likely that observations are obtained from every component.
MCMCs were run for 30,000 iterations, with the first 20,000 discarded as burn-in. We assume that $M-1\mid\Lambda\sim\text{Poisson}(\Lambda)$ and $\Lambda\sim\text{Gamma}(1,0.1)$ and employ a multivariate normal kernel with a normal–inverse-Wishart hierarchical prior in Section 3.4.1, where $b_{0}=\mathrm{median(y)},B_{0}=\mathrm{diag}(R_{1}^{2},\dots,R_{d}^{2})$, $R_{j}$ is the range of data $y_{1:n}$ in $j$th dimension, $c_{0}=2.5+(d-1)/2,g_{0}=0.5+(r-1)/2$ and $G_{0}=100g_{0}/c_{0}\mathrm{diag}(1/R_{1}^{2},\dots,1/R_{d}^{2})$. Also, we set shape parameters $\alpha,\gamma=1.0$ for MFM-IGau, DMFM-IGau and MFM-Ga, and we assume a prior F distribution $F(6,3)$ to shape parameter DMFM-Ga as in fruhwirth2021generalized.
To measure performance, we consider the posterior mean of the Adjusted Rand index (ARI) (hubert1985comparing), which is a measure of the goodness of clustering. When it is close to $1$, the assignment estimate is reasonable. The respective averages over $50$ repetitions are denoted by $\widehat{\mathrm{ARI}}$.
Overall, figure 2 shows that $\widehat{\mathrm{ARI}}$ of all methods increase toward $1$ as $n$ becomes larger. MFM-IGau consistently attains higher clustering accuracy than MFM-Ga, and DMFM-IGau exhibits the best overall performance. Although the difference between DMFM-IGau and DMFM-Ga is small, DMFM-IGau does not incorporate learning for $\alpha$. This result indicates that stable posterior inference is achieved due to the less informative nature of the normalized inverse Gaussian distribution.
Figure 2: Line plots of the posterior mean of ARI averaged over $50$ Monte Carlo replications for MFM-Ga, DMFM-Ga, MFM-IGau and DMFM-IGau.